4.) What volume of oxygen gas is produced when 19.6 g of mercury(II) oxide reacts completely according to the following reaction at 25 °C and 1 atm?

mercury(II) oxide (s) mercury (l) + oxygen(g)

_____liters oxygen gas

5.) A mixture of nitrogen and argon gases, in a 7.75 L flask at 21 °C, contains 7.83 grams of nitrogen and 12.4 grams of argon. The partial pressure of argon in the flask is atm and the total pressure in the flask is _____ atm.

6.) A mixture of carbon dioxide and oxygen gases is maintained in a 5.26 L flask at a pressure of 3.14 atm and a temperature of 84 °C. If the gas mixture contains 16.9 grams of carbon dioxide, the number of grams of oxygen in the mixture is ____g.

7.) A mixture of methane and argon gases, at a total pressure of 798 mm Hg, contains 2.77 grams of methane and 3.40 grams of argon. What is the partial pressure of each gas in the mixture?
PCH4 = ____mm Hg
PAr = ____ mm Hg
1
Expert's answer
2020-10-16T08:57:45-0400

4. ( (273)(1)) =19.6

((298)(1) = ?

= ((19.6)(298))/ 273


= 5840.8/273

= 21.39g

5. 12.4g- 7.83g

= 4.57g


6. ((3.14)(571))/ 273


= 1792.94/273

= 6.57g


7. ((798mmhg)(3.40g))/ 2.77g


= 27132/2.77


= 979.49mmHg